OpenStudy (anonymous):

The population in the world in 2006 was 6.53 billion people and was growing at a rate of 1.14% per year. Assuming that this growth rate continues, the model represents the population P (in billions of people) in year t. Answer the following questions. [A] What will be the anticipated world population in 2020? [B] For what value of t will the world population be 12 billion? What date corresponds to this value of t?

OpenStudy (anonymous):

nobody know the answer to this one

OpenStudy (akshay_budhkar):

this is a question on differential equations...:) are you sure it is there for you?

OpenStudy (anonymous):

what do you mean?

OpenStudy (anonymous):

yes we can do this

OpenStudy (anonymous):

@akshay did you get your probability answer?

OpenStudy (anonymous):

@treyhud the formula is \[P(t)=6.53e^{.0114t}\] where t is years after 2006

OpenStudy (anonymous):

2020 is 14 years after 2006 to replace t by 14 to get answer to first part

OpenStudy (akshay_budhkar):

yes.. satellite... but its good you r helping.. i am not in touch with it for 2 years.. thanks:) i was searching for the format.

OpenStudy (anonymous):

\[.0114\times 14=.1596\] and \[6.53e^{.1596}=7.64\] rounded

OpenStudy (anonymous):

@akshay look at the very end of that problem gives and easy way to do it without using a solid. very simple actually.

OpenStudy (anonymous):

we can work through it if you like

OpenStudy (anonymous):

P(t)=6.53(1.8377)^r-2006=12 so the value of t is 1.8377 ? i dont know if that is right

OpenStudy (anonymous):

@treyhud answer to second question put \[12=6.53e^{.0114t}\] and solved for t

OpenStudy (anonymous):

t is number of years after 2006

OpenStudy (akshay_budhkar):

@satellite i am not getting you??

OpenStudy (anonymous):

for this problem or probability one?

OpenStudy (akshay_budhkar):

probability one..

OpenStudy (anonymous):

ok let me finish this one and then start another thread

OpenStudy (akshay_budhkar):

ok sure.

OpenStudy (anonymous):

@treyhud is it clear what is going on here? we do not work with the numbers 2006,2020 etc. instead you put 2006 = 0 that is when we start counting 2020=14 that is 14 years later

OpenStudy (anonymous):

ok yeah that makes since with the years part. I am just trying to solve for t for the second part. What are you getting cause my number seems a little off

OpenStudy (anonymous):

formula for continuous grown it \[P=P_0e^{rt}\] where \[P_0\] is initial population, r is rate as a decimal and t is time

OpenStudy (anonymous):

in the last problem you have P=12 so put \[12=6.53e^{.0114t}\] and solve for t

OpenStudy (anonymous):

you do this via \[\frac{12}{6.53}=e^{.0114t}\] \[\ln(\frac{12}{6.53})=.0114t\] \[t=\ln(\frac{12}{6.53})\div .0014\]

OpenStudy (anonymous):

so it would look like this \[t=6.53e^.0114/12\]

OpenStudy (anonymous):

oh no. you have to get the variable out of the exponent by taking the log

OpenStudy (anonymous):

ok thats my problem then

OpenStudy (anonymous):

first step is divide both sides by 6.53 second step is take the natural log of both sides

OpenStudy (anonymous):

then divide the result by .0014

OpenStudy (anonymous):

you get \[t=\ln(\frac{12}{6.53})\div .0014\]

OpenStudy (anonymous):

434.64?

OpenStudy (anonymous):

i made a typo sorry. you need to divide by .0114, not .0014

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

that is the rate of growth, .0114

OpenStudy (anonymous):

should have written \[t=\ln(\frac{12}{6.53})\div .0114\]

OpenStudy (anonymous):

my mistake. i copied wrong. you get around 53

OpenStudy (anonymous):

so 53 years after 2006 is 2059

OpenStudy (anonymous):

gotcha. thanks alot. u taught me better on this chat deal than my professor has all semester

OpenStudy (anonymous):

i will take that as a compliment. yw